(10 pts) A firm has the following short run total costs, where Q is output and...

  1. (10 pts) A firm has the following short run total costs, where Q is output and TC is total cost:

Q

TC

0

$ 100

1

110

2

130

3

160

4

200

5

250

6

310

7

380

8

460

9

550

10

650

11

760

  1. What is total fixed cost equal to?
  1. What is average total cost at Q = 3?

  1. What is average variable cost at Q = 6?

  1. What is marginal cost at Q = 8?

  1. At Q=8, is the firm operating under increasing or decreasing returns? Why?
Answer #1

a.

Total fixed cost is that cost which is born when there is zero unit of output produced.

Hence TFC=$100

b.

When Q=3

ATC=TC\Q

=160\3

=53.33

c.

When Q=6

TVC=TC-TFC

=310-100

=210

AVC=TVC\Q

=210\6

=35

d.

MC=TCn-TCn-1

when Q=8

MC=TC8- TC7

=460-380

=80